OpenStudy (anonymous):

Circle S has an equation of (x + 16)^2 + (y – 9)^2 = 4. What is the center and radius of circle S? Answer Center: (16, -9); Radius: 4 Center: (-16, 9); Radius: 4 Center: (16, -9); Radius: 2 Center: (-16, 9); Radius: 2

OpenStudy (anonymous):

do you know the formula of the circle ?

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] where (h,k) is the center and r is the radius

OpenStudy (anonymous):

want to take a stab at the answer, tahtah?

OpenStudy (amistre64):

thats why i just call h Cx and k Cy

OpenStudy (anonymous):

uhhhhhhmmmmmmmm one sec

OpenStudy (anonymous):

you have it in just the form you want \[(x + 16)^2 + (y – 9)^2 = 4\] \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

identify h, k and r

OpenStudy (anonymous):

would it be A?

OpenStudy (anonymous):

definitely not a, because r^2=4 r=2

OpenStudy (anonymous):

Oh im slow. So is it C then?

OpenStudy (anonymous):

@Dockworker \[\Large if \quad r^2=4 \longrightarrow r=\pm\sqrt{4} \quad ??\]

OpenStudy (anonymous):

not c either, its d you can rewrite the equation to better help you see \[(x-(-16))^2+(y-9)^2=2^2\]

OpenStudy (anonymous):

@Kreshnik, we're dealing with a radius here. you can discard the negative value.

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

oh darn i was not close at all. Thank you <3

OpenStudy (anonymous):

\[r^2=4\implies r=2\] \[16=-h\implies h=-16\] \[-9=-k\implies k=9\] center is \((-16,9)\) radius is 2